Calculate the pH of the solution to see if it is in an acceptable range (pH between 6 and 8). If not, please suggest what to add (you do not have to calculate how much) to get within the acceptable range. The current volume and concentrations of acids and bases are as follows:

13.57 mL of 0.007213 M NH4OH

2.39 L of 0.01053 M C5H5N

11.56 L of 0.0026567 M N(CH3)3

13.27 mL of 0.086031 M NH3

5.23 L of 0.3119 M CH3NH2

813.14 mL of 0.005629 M glycine

3.53 L of 0.003501 M H3PO4

45.6 mL of 0.05304 M CH3COOH

2.15 L of 0.1234 M H3C6H5O7

348.29 mL of 0.02150 M C6H8O7

896.25 mL of 0.3587 M H2C2O4

1.29 mL of 0.07389 M HNO2

37.29 mL of 0.02319 M disodium EDTA

2.71 mL of 18.02 M NaOH

3.59 L of 0.1649 M HCl
Answer #1

0 H20 cene, of nittyolt a 0.007213 M. Nl4got I rita + oH] It produces th] Hoody [ot] = 0.00723 M as it is of audity 1 рон е и[ests n] = [of] = 0.01053 m Han pH = 14-pot - 14+ log 0.010s3 and pH = 12.02 So the solution is basic and to you need to and6 NHg is also hasiz. s stilong with v erot NH3 + H20 T NH4 + oH so [Nta] = [OH) = 0, 086031 M... Bolo la lokasitas pH = 14-po